# maximum a posteriori

### v1

@@ 4 Experiments and Results
To establish the benefits of the lipchitz BOLT-GAN design, we conduct experiments on standard image generation tasks using the MNIST (LeCun et al.,, 1998) dataset. and CelebA (Liu et al.,, 2015). We use the Fréchet inception distance (FID) (Heusel et al.,, 2017) to assess the quality of the generated images over training epochs and provide comparisons to the WGAN.
@@ 4.1 Experimental Setup
We adopt a residual DCGAN backbone (Gulrajani et al.,, 2017): the generator upsamples a set of 128-D Gaussian latent images to $64{\times}64$ RGB via residual blocks; the critic symmetrically downsamples these images with global-sum pooling and a final linear head for classification. BatchNorm/ReLU activation is used in the generator. LeakyReLU in the critic, with a Lipschitz constraint enforced through a gradient penalty ($\lambda_{GD}=10$). We evaluate performance on CIFAR-10, CelebA-64, and LSUN Bedroom/Church-64. For training we use Adam ($\alpha=2{\times}10^{-4},\beta_{1}=0.5,\beta_{2}=0.999$), with batch size 64, for 20 epochs. FID (Heusel et al.,, 2017) is computed every 5 epochs on 10,000 samples using torch-fidelity (Seitzer,, 2020), averaged over three seeds. Additional architectural and hardware details appear in the supplementary.
4.2 Quantitative and Convergence Results
Table 3 shows that BOLT–GAN consistently surpasses WGAN, achieving 10–60% lower FID across all benchmarks. As illustrated in Figure 1, it also converges faster and exhibits smoother training dynamics with reduced variance and fewer oscillations. Complementary qualitative results in Figure 2 highlight sharper textures and fewer artifacts in samples generated by BOLT–GAN after 100 epochs on CIFAR-10. We also trained BOLT–GAN without Lipschitz regularization. This TV-based variant quickly diverges (FID >>300–400), confirming that the gradient-penalty term is essential for stability.
Table 1: FID (↓) after 20 epochs for the non-Lipschitz BOLT–GAN ($\lambda_{\mathrm{GP}}=0$) compared with the Lipschitz version.
Dataset Non-Lipschitz ($\lambda_{\mathrm{GP}}=0$) Lipschitz ($\lambda_{\mathrm{GP}}=10$)
CIFAR-10 $\approx$350 (unstable) 44.2 ± 1.2
CelebA-64 $\approx$300 9.2 ± 0.4
LSUN Bedroom-64 Diverged (>>400) 40.3 ± 2.3
4.3 Ablation Studies
@@
Table 2: FID (↓) of WGAN-GP and BOLT-GAN-GP on CIFAR-10 across different values of $\lambda_{\mathrm{GP}}$.
Method $\lambda_{\mathrm{GP}}=1$ 5 10 20
WGAN-GP 60.8 59.6 60.0 61.1
BOLT-GAN-GP 44.3 43.6 44.2 46.7
Table 3: FID ($\downarrow$) after 20 epochs.
Dataset WGAN BOLT GAN Relative $\Delta$ (%)
CIFAR-10 $60.0\!\pm\!1.8$ $44.2\!\pm\!1.2$ 26.3-26.3
CelebA-64 $10.3\!\pm\!0.5$ $9.2\!\pm\!0.4$ 10.7-10.7
LSUN Bedroom-64 $102.5\!\pm\!2.9$ $40.3\!\pm\!2.3$ 60.7-60.7
LSUN Church-64 $43.5\!\pm\!1.4$ $14.8\!\pm\!0.6$ 65.9-65.9
Table 4: Best FID ($\downarrow$) after 20 epochs for different $\lambda_{\text{GP}}$. Bold marks the default used in all other experiments.
Dataset $\lambda_{\text{GP}}$
1 5 10 20
[…]
CelebA-64 $10.5\!\pm\!0.5$ $9.4\!\pm\!0.4$ $\mathbf{9.2\!\pm\!0.4}$ $10.0\!\pm\!0.5$
Table 5: CIFAR‑10 FID at 20 ep with balanced vs. imbalanced prior.
Configuration FID $\Delta$ vs. $\pi=0.50$
Balanced ($\pi=0.50$) $44.2$
Imbalanced ($\pi=0.45$) $\mathbf{43.1}$  $2.4\%$
020204040606080801001005050100100150150EpochFID (lower is better)WGAN (mean & 95% CI)BOLT GAN(mean & 95% CI)
[…]
@@ 5 Conclusion
This paper has presented BOLT–GAN, a streamlined modification of the WGAN framework inspired by the Bayes optimal learning threshold (BOLT). Through extensive experiments on four diverse image‐generation benchmarks, namely CIFAR-10, CelebA-64, LSUN Bedroom-64, and LSUN Church‐64, we have demonstrated that BOLT–GAN consistently produces smoother training dynamics and lowers FID scores by 10-60% compared to the standard WGAN. These gains hold across both low‐ and high‐resolution settings, underscoring the robustness and generality of our approach. The current study introduces two extreme cases of BOLT-GAN, namely the vanilla and Lipschitz BOLT-GAN. While the former minimizes the total variation, the latter restricts a metric that is weaker than the Wasserstein distance. Exploring constraints under which BOLT-GAN framework restricts alternative measures remains an interesting direction for future research.
@@ A.1 Metrics
Lemma 2 (Equivalent forms of total variation (Dudley, 2002b, )).
@@
Lemma 3 (Kantorovich–Rubinstein duality for $W_{1}$(Villani et al.,, 2009)).
@@
Lemma 4 (Bounding $W_{1}$ by TV on bounded spaces (Villani et al.,, 2009)).
@@ A.2 BOLT and BOLT–GAN
For a measurable $h:\mathcal{X}\to[-1,0]$,
$\mathcal{L}_{\mathrm{BOLT}}(h)\;=\;q_{2}-\Big(q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)]-q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]\Big).$ (57)
The universal bound of Naeini et al. (2025, Thm. 1) implies that for any such $h$, $\varepsilon_{\mathrm{Bayes}}\leq\mathcal{L}_{\mathrm{BOLT}}(h)$.
Risk and per-example BOLT loss
Let $Y\in\{1,2\}$, $X\mid(Y=i)\sim P_{i}$, and let $h:\mathcal{X}\to[-1,0]$ be measurable. Define $s(y)\vcentcolon=\mathbbm{1}\{y=2\}-\mathbbm{1}\{y=1\}\in\{-1,+1\}$ (or equivalently, $s(y)=(-1)^{y}$), . We further define the per-example BOLT loss and the population risk:
$\ell_{\mathrm{BOLT}}(z,y)$ $\;\vcentcolon=\;s(y)\,z,$ (58)
$R(h)$ $\;\vcentcolon=\;\mathbb{E}\!\big[\ell_{\mathrm{BOLT}}(h(X),Y)\big]\;=\;-\,q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)]\;+\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)].$ (59)
With the BOLT functional from Def. 5,
$\mathcal{L}_{\mathrm{BOLT}}(h)$ $=q_{2}\;-\;\Big(q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)]\;-\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]\Big)$ (60)
$=q_{2}\;+\;R(h),$ (61)
so the minimizers of $R(h)$ coincide with the minimizers of $\mathcal{L}_{\mathrm{BOLT}}(h)$.
@@ Appendix B Proofs and Further Results on BOLT
Theorem 5 (BOLT Estimator of MAP Classifier (Theorem 2 from main paper)).
Let $(X,Y)$ describe a binary classification problem with $Y\in\{1,2\}$, priors $q_{1}=\Pr\{Y=1\}$ and $q_{2}=\Pr\{Y=2\}$, class‑conditional laws $P_{1}$ and $P_{2}$, and posterior
$\eta(x)\;\vcentcolon=\;\Pr\{Y=1\mid X=x\}.$ (64)
Let $h^{\star}\in\arg\min_{h:\,\mathcal{X}\to[-1,0]}\mathcal{L}_{\mathrm{BOLT}}(h)$ (or equivalently, $h^{\star}\in\arg\min R(h)$). Then, $h^{\star}$ is pointwise optimal almost surely and takes only the endpoint values:
$h^{\star}(x)\;=\;\begin{cases}0,&\eta(x)>\tfrac{1}{2},\\[2.0pt] -1,&\eta(x)<\tfrac{1}{2},\\[2.0pt] \text{any }z\in[-1,0],&\eta(x)=\tfrac{1}{2}.\end{cases}$ (65)
Consequently, the plug‑in classifier at threshold 0.5-0.5, i.e.
$\widehat{C}(x)\;=\;1\{\,h^{\star}(x)\geq-0.5\,\},$ (66)
coincides almost everywhere with the MAP decision rule $C_{\mathrm{MAP}}(x)=1\{\eta(x)\geq\tfrac{1}{2}\}$. Moreover, if a model $h_{\theta}$ satisfies $\|h_{\theta}-h^{\star}\|_{\infty}\leq\epsilon$, then the plug‑in decision
$\widehat{C}_{\theta}(x)\;=\;\begin{cases}C_{1},&h_{\theta}(x)\geq-0.5,\\[2.0pt] C_{2},&h_{\theta}(x)<-0.5,\end{cases}$ (67)
is a plug‑in estimator of the MAP classifier.
@@
Fix $x\in\mathcal{X}$ and abbreviate $\eta=\eta(x)$. The conditional risk of predicting a value $z\in[-1,0]$ is
$r_{\eta}(z)$ $\;\vcentcolon=\;\mathbb{E}\!\big[\ell_{\mathrm{BOLT}}(z,Y)\mid X=x\big]\;$ (68)
$=\;\eta\,\ell_{\mathrm{BOLT}}(z,1)+(1-\eta)\,\ell_{\mathrm{BOLT}}(z,2)=$ (69)
$\;\eta(-z)+(1-\eta)z\;=\;(1-2\eta)\,z.$ (70)
To minimize $r_{\eta}(z)$ over $z\in[-1,0]$, we can say
  1. 1.
    If $\eta>\tfrac{1}{2}$, then $1-2\eta<0$. Thus, $r_{\eta}(z)$ is strictly decreasing in $z$. The minimum over [1,0][-1,0] in this case is attained at the maximal $z=0$.
  2. 2.
    If $\eta<\tfrac{1}{2}$, then $1-2\eta>0$. Thus, $r_{\eta}(z)$ is strictly increasing in $z$. The minimum therefore occurs at the minimal z=1z=-1.
  3. 3.
    If $\eta=\tfrac{1}{2}$, then $r_{\eta}(z)\equiv 0$ on [1,0][-1,0], and hence any $z\in[-1,0]$ is optimal.
This yields the pointwise characterization (65). Since $\widehat{C}(x)=1\{h^{\star}(x)\geq-0.5\}$ equals $1\{\eta(x)\geq\tfrac{1}{2}\}$ by (65), the plug‑in rule (66) coincides with MAP almost everywhere. The final claim for $h_{\theta}$ follows the fact that $\|h_{\theta}-h^{\star}\|_{\infty}\leq\epsilon$ leaves the decision boundary at 0.5-0.5 unchanged whenever $h^{\star}\in\{-1,0\}$. ∎
Remark 4 (Alternative normalization).
If one parametrizes $h:\mathcal{X}\to[0,1]$ (e.g., the bounded score after sigmoid activation), the same argument applied to the loss $\tilde{\ell}(z,y)=(-1)^{y}(2z-1)$ yields $h^{\star}(x)=1\{\eta(x)>\tfrac{1}{2}\}$ and a threshold at +0.5+0.5. The two normalizations are related by an affine map, which adjusts the threshold accordingly.
Corollary 2 (Bayes consistency of the plug‑in classifier).
Assume $\Pr\!\big(\eta(X)=\tfrac{1}{2}\big)=0$ under the data distribution of $(X,Y)$. Let $\widehat{h}_{n}:\mathcal{X}\to[-1,0]$ be risk consistent for $\ell_{\mathrm{BOLT}}$, i.e.
$R(\widehat{h}_{n})\;\xrightarrow[n\to\infty]{}\;\inf_{h}R(h)\;=\;R(h^{\star}),$ (71)
with $R$ as in (59). Define $\widehat{C}_{n}(x)=1\{\widehat{h}_{n}(x)\geq-0.5\}$. Then the 011 risk converges to the Bayes risk:
$\Pr\!\big(\widehat{C}_{n}(X)\neq Y\big)\;\xrightarrow[n\to\infty]{}\;\Pr\!\big(C_{\mathrm{MAP}}(X)\neq Y\big).$ (72)
@@
Write the conditional excess risk at $x$ as
$r_{\eta(x)}\big(\widehat{h}_{n}(x)\big)-r_{\eta(x)}\big(h^{\star}(x)\big)\;=\;|1-2\eta(x)|\cdot\big|\,\widehat{h}_{n}(x)-h^{\star}(x)\,\big|.$ (73)
Then, by (71), we can write
$R(\widehat{h}_{n})-R(h^{\star})=\mathbb{E}\!\big[|1-2\eta(X)|\,|\widehat{h}_{n}(X)-h^{\star}(X)|\big]\to 0.$ (74)
Fix any $\delta\in(0,1)$ and split the misclassification event as follows:
$\mathcal{E}_{n}\;\vcentcolon=\;\{\widehat{C}_{n}(X)\neq C_{\mathrm{MAP}}(X)\}\;=\;\big(\mathcal{E}_{n}\cap\{|1-2\eta(X)|\geq\delta\}\big)\;\cup\;\{|1-2\eta(X)|<\delta\}.$ (75)
For the case $|1-2\eta(X)|\geq\delta$, a decision mismatch implies $|\widehat{h}_{n}(X)-h^{\star}(X)|\geq\tfrac{1}{2}$ because $h^{\star}\in\{-1,0\}$ and the plug‑in threshold is 0.5-0.5. Therefore,
$R(\widehat{h}_{n})-R(h^{\star})$ $\;\geq\;\mathbb{E}\!\big[\,|1-2\eta(X)|\,\tfrac{1}{2}\,\mathbf{1}_{\mathcal{E}_{n}\cap\{|1-2\eta(X)|\geq\delta\}}\,\big]$ (76)
$\;\geq\;\tfrac{\delta}{2}\,\Pr\!\big(\mathcal{E}_{n}\cap\{|1-2\eta(X)|\geq\delta\}\big).$ (77)
Letting $n\to\infty$ gives $\Pr(\mathcal{E}_{n}\cap\{|1-2\eta(X)|\geq\delta\})\to 0$. Moreover, by assumption, $\Pr(|1-2\eta(X)|<\delta)\to 0$ as $\delta\downarrow 0$. We hence first send $n\to\infty$, and then $\delta\downarrow 0$ in (75) to obtain (72). ∎

### v2

@@ 5 Experiments and Results
To evaluate the proposed BOLT-GAN, we conduct unconditional image-generation experiments on four standard benchmarks: CIFAR-10, CelebA-64, LSUN Bedroom-64, and LSUN Church-64. We compare against (i) WGAN, since it uses the same 1-Lipschitz enforcement mechanism as our implementation, and (ii) a HINGE-GAN, a widely used stability-focused baseline. Unless otherwise stated, BOLT-GAN and WGAN use the same generator and discriminator architectures and optimizer settings, i.e., the only difference is the training objective. In the experiments, we report the Fréchet inception distance (FID), as well as Improved Precision and Recall (Kynkäänniemi et al., 2019) to quantify sample fidelity and mode coverage.
@@ 5.1 Experimental Setup
020204040606080801001005050100100150150EpochFID (lower is better)WGAN-GP (mean & 95% CI)BOLT-GAN(mean & 95% CI)
Figure 1: CIFAR-10 FID vs. epochs (mean and 95% confidence interval over 3 seeds), showing consistently improved convergence of BOLT-GAN relative to WGAN-GP.
@@
5.2 Quantitative Results
Figure 1 shows the CIFAR-10 FID trajectory over training, where BOLT-GAN consistently achieves lower FID than WGAN-GP across epochs. Table 1 reports FID after 20 epochs. BOLT-GAN improves substantially over WGAN across all datasets and is competitive with (and in several cases slightly better than) HINGE-GAN under the same compute budget. To quantify mode coverage, Table 2 reports Precision and Recall (higher is better). Across datasets where we computed the metric, BOLT-GAN improves both precision and recall relative to WGAN, indicating better quality and coverage. Figure 2 shows CIFAR-10 samples after 100 epochs, where BOLT-GAN yields sharper images with fewer artifacts compared to WGAN.
[…]
Dataset HINGE-GAN BOLT-GAN WGAN
CIFAR-10 $6.9\pm 0.8 44$.2 $\pm$ 0.6 $60.0\pm 1.0$
LSUN Church-64 $19.7\pm 0.5$ 14.8 $\pm$ 0.4 $43.5\pm 0.9$
LSUN Bedroom-64 $28.8\pm 0.6$ 28.4 $\pm$ 0.5 $102.5\pm 2.0$
CelebA-64 $19.6\pm 0.5$ 9.2 $\pm$ 0.3 $10.3\pm 0.4$
Table 2: Improved Precision/Recall ($\uparrow$) with $n=2000$, $k=3$ (higher is better). Best values per dataset/epoch are bolded.
Dataset Epochs Method Precision Recall
CIFAR-10 20 BOLT-GAN 0.763 0.352
20 WGAN 0.714 0.183
100 BOLT-GAN 0.812 0.469
100 WGAN 0.764 0.102
LSUN Church-64 20 BOLT-GAN 0.750 0.411
20 WGAN 0.681 0.176
100 BOLT-GAN 0.792 0.502
100 WGAN 0.731 0.098
LSUN Bedroom-64 20 BOLT-GAN 0.598 0.450
20 WGAN 0.388 0.098
CelebA-64 20 BOLT-GAN 0.840 0.548
20 WGAN 0.672 0.104
Table 3 compares Lipschitz discriminator ($\lambda_{\rm GP}=10$) with an unconstrained one ($\lambda_{\rm GP}=0$). The results show that removing Lipschitz enforcement, i.e., setting $\lambda_{\rm{GP}}=0$, leads to severe numerical instability, while training via a Lipschitz-constrained discriminator remains stable.
Table 3: FID ($\downarrow$) after 20 epochs for Max-BER (unregularized; $\lambda_{\rm GP}=0$) compared with the proposed BOLT-GAN (Lipschitz; $\lambda_{\rm GP}=10$).
Dataset   Max-BER BOLT-GAN
CIFAR-10 $352\pm 48$ $44.2\pm 1.2$
CelebA-64 $276\pm 35$ $9.2\pm 0.4$
LSUN Church-64 $468\pm 56$ $31.6\pm 0.7$
5.3 Ablation Studies
Label smoothing can be accomplished in BOLT-GAN by changing the prior $\pi$. Table 4 shows the FID at 20 and 100 epochs for various choices of prior $\pi$. As the results show, a modest prior imbalance can slightly improve FID on CIFAR-10.
[…]
Prior $\pi$ FID @ 20 epochs FID @ 100 epochs
$0.50$ $44.2\pm 0.6$ $36.6\pm 0.8$
$0.45$ $45.3\pm 0.7$ $\mathbf{36.2\pm 0.7}$
$0.40$ $48.5\pm 1.0$ $41.0\pm 1.1$
$0.35$ $51.4\pm 1.3$ $44.8\pm 1.4$
Table 5: FID ($\downarrow$) of WGAN and BOLT-GAN on CIFAR-10 across different values of $\lambda_{\rm GP}$.
$\lambda_{\rm GP}$ 1 5 10 20
WGAN 60.8 59.6 60.0 61.1
BOLT-GAN 44.3 43.6 44.2 46.7
We next test the robustness to gradient-penalty strength, we sweep $\lambda_{\rm GP}\in\{1,5,10,20\}$ for 20 epochs and report FID averaged over N seeds; see Table 5. The results show that performance stays within roughly $\pm 1$ FID across the range. Figure 3 further reports FID trajectories across several GP strengths, showing only mild sensitivity over $\lambda_{\rm GP}\in\{1,5,10,20\}$ and motivating our default choice $\lambda_{\rm GP}=10$.
@@ 6 Conclusion
We proposed BOLT-GAN, a Bayes-error-motivated objective for stable GAN training that trains a bounded 1-Lipschitz discriminator with the BOLT loss. The discriminator in this case estimates a Bayes optimal binary classifier for the real-fake classification task, and the generator learns the data distribution adversarially by maximizing the discrimination BER. We showed that the Max-BER problem solved by BOLT-GAN minimized an IPM between the generator and data distributions that is upper-bounded by the Wasserstein-1 distance. In absence of Lipschitz constraint at the discriminator, the Max-BER problem minimizes the TV between the two distributions, leading to empirical instability. Experiments demonstrate that BOLT-GAN improves FID and Improved Precision/Recall over the state-of-the-art frameworks, such as WGAN and HINGE-GAN, under identical architectures and training budgets. Understanding how alternative constraints at the discriminator impact the convergence behavior of the Max-BER approach for training is an interesting direction for future research.
@@ A.1 Metrics
Lemma 3 (Equivalent forms of total variation (Dudley, 2002b)).
@@
Lemma 4 (Kantorovich–Rubinstein duality for $W_{1}$(Villani and others, 2009)).
@@
Lemma 5 (Bounding $W_{1}$ by TV on bounded spaces (Villani and others, 2009)).
@@ A.2 BOLT and BOLT–GAN
For a measurable $h:\mathcal{X}\to[0,1]$, define
$\mathcal{L}_{\mathrm{BOLT}}(h)\;\coloneqq\;q_{1}\;+\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]\;-\;q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)].$ (57)
The universal bound implies that for any such $h$,
$\varepsilon_{\rm Bayes}\;\leq\;\mathcal{L}_{\mathrm{BOLT}}(h).$ (58)
Risk and per-example BOLT loss.
Let $Y\in\{1,2\}$, $X\mid(Y=i)\sim P_{i}$, and let $h:\mathcal{X}\to[0,1]$ be measurable. Define
$s(y)\coloneqq\mathbf{1}\{y=2\}-\mathbf{1}\{y=1\}\in\{-1,+1\},$ (59)
and the per-example loss and population risk by
$\ell_{\rm BOLT}(z,y)$ $\coloneqq s(y)\,z,$ (60)
$R(h)$ $\coloneqq\mathbb{E}\!\left[\ell_{\rm BOLT}(h(X),Y)\right]=-q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)]+q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)].$ (61)
With (57),
$\mathcal{L}_{\mathrm{BOLT}}(h)=q_{1}+R(h),$ (62)
so minimizers of $R(h)$ coincide with minimizers of $\mathcal{L}_{\mathrm{BOLT}}(h)$.
@@ Appendix B Proofs and Further Results on BOLT
Theorem 6 (BOLT estimator of MAP classifier).
Let $(X,Y)$ describe a binary classification problem with $Y\in\{1,2\}$, priors $q_{1}=\Pr\{Y=1\}$ and $q_{2}=\Pr\{Y=2\}=1-q_{1}$, class-conditional laws $X\mid(Y=i)\sim P_{i}$, and posterior
$\eta(x)\coloneqq\Pr\{Y=1\mid X=x\}.$ (65)
Consider the BOLT upper bound in the $[0,1]$ normalization (as in the main paper): for measurable $h:\mathcal{X}\to[0,1]$ define
$\mathcal{B}(h)\;\coloneqq\;q_{1}\;+\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]\;-\;q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)].$ (66)
Let $h^{\star}\in\arg\min_{h:\mathcal{X}\to[0,1]}\mathcal{B}(h)$. Then $h^{\star}$ is pointwise optimal almost surely and satisfies, for $(P_{1}+P_{2})$-a.e. $x$,
$h^{\star}(x)=\begin{cases}1,&\eta(x)>\tfrac{1}{2},\\ 0,&\eta(x)<\tfrac{1}{2},\\ \text{any }z\in[0,1],&\eta(x)=\tfrac{1}{2}.\end{cases}$ (67)
Consequently, the plug-in classifier at threshold $1/2$,
$\widehat{C}(x)\;=\;\mathbf{1}\!\left\{h^{\star}(x)\geq\tfrac{1}{2}\right\},$ (68)
coincides almost everywhere with the MAP decision rule $\mathbf{1}\{\eta(x)\geq\tfrac{1}{2}\}$.
Moreover, if a learned model $h_{\theta}$ satisfies $\|h_{\theta}-h^{\star}\|_{\infty}\leq\varepsilon$ for some $\varepsilon<\tfrac{1}{2}$, then the plug-in decision
$\widehat{C}_{\theta}(x)\;=\;\mathbf{1}\!\left\{h_{\theta}(x)\geq\tfrac{1}{2}\right\}$ (69)
agrees with $\widehat{C}(x)$ (and hence with MAP) except possibly on the decision boundary set $\{x:\eta(x)=\tfrac{1}{2}\}$.
@@
The additive constant $q_{1}$ does not affect minimizers, so minimizing $\mathcal{B}(h)$ over $h$ is equivalent to minimizing
$R(h)\;\coloneqq\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]\;-\;q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)].$ (70)
Fix $x\in\mathcal{X}$ and abbreviate $\eta=\eta(x)$. The conditional risk of predicting a value $z\in[0,1]$ is
$r_{\eta}(z)$ $\coloneqq\mathbb{E}\!\left[\,\ell(z,Y)\mid X=x\,\right]$ (71)
$=\eta\,\ell(z,1)+(1-\eta)\,\ell(z,2),$
where (consistent with the main text) $\ell(z,1)=-z$ and $\ell(z,2)=+z$. Hence
$r_{\eta}(z)$ $=\eta(-z)+(1-\eta)z$
$=(1-2\eta)\,z.$ (72)
To minimize $r_{\eta}(z)$ over $z\in[0,1]$:
  • If $\eta>\tfrac{1}{2}$ then $1-2\eta<0$ and $r_{\eta}(z)$ is decreasing in $z$, so the minimum is at $z=1$.
  • If $\eta<\tfrac{1}{2}$ then $1-2\eta>0$ and $r_{\eta}(z)$ is increasing in $z$, so the minimum is at $z=0$.
  • If $\eta=\tfrac{1}{2}$ then $r_{\eta}(z)\equiv 0$ and any $z\in[0,1]$ is optimal.
This yields the pointwise characterization (67).
For the plug-in rule, when $\eta(x)\neq\tfrac{1}{2}$ we have $h^{\star}(x)\in\{0,1\}$, and
$h^{\star}(x)\geq\tfrac{1}{2}\;\Longleftrightarrow\;h^{\star}(x)=1\;\Longleftrightarrow\;\eta(x)>\tfrac{1}{2},$ (73)
so (68) agrees with MAP almost everywhere (ties only when $\eta(x)=\tfrac{1}{2}$).
Finally, if $\|h_{\theta}-h^{\star}\|_{\infty}\leq\varepsilon<\tfrac{1}{2}$ and $h^{\star}(x)\in\{0,1\}$, then $h_{\theta}(x)$ cannot cross the threshold $1/2$, so (69) matches (68) away from the boundary set $\{\eta=\tfrac{1}{2}\}$. ∎
Corollary 3 (Bayes consistency of the plug-in classifier).
Assume
$\Pr\!\left(\eta(X)=\tfrac{1}{2}\right)=0,$ (74)
under the distribution of $(X,Y)$, where $\eta(x)=\Pr\{Y=1\mid X=x\}$. Let $\widehat{h}_{n}:\mathcal{X}\to[0,1]$ be risk-consistent for the BOLT loss $\ell(z,1)=-z$ and $\ell(z,2)=+z$, i.e.
$R(\widehat{h}_{n})\xrightarrow[n\to\infty]{}\inf_{h:\mathcal{X}\to[0,1]}R(h)=R(h^{\star}),$ (75)
where
$R(h)\coloneqq\mathbb{E}\!\left[\ell\!\left(h(X),Y\right)\right]=-q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)]+q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)].$ (76)
Define the plug-in classifier at threshold $1/2$,
$\widehat{C}_{n}(x)\coloneqq\mathbf{1}\!\left\{\widehat{h}_{n}(x)\geq\tfrac{1}{2}\right\}.$ (77)
Then the 011 risk converges to the Bayes risk:
$\Pr\!\left(\widehat{C}_{n}(X)\neq Y\right)\xrightarrow[n\to\infty]{}\Pr\!\left(C_{\rm MAP}(X)\neq Y\right).$ (78)
@@
Let $h^{\star}$ be a population minimizer from Theorem 6 (your $[0,1]$ version), so $h^{\star}(x)\in\{0,1\}$ for $\eta(x)\neq\tfrac{1}{2}$. Fix $x\in\mathcal{X}$ and abbreviate $\eta=\eta(x)$. The conditional risk of predicting $z\in[0,1]$ is
$r_{\eta}(z)\coloneqq\mathbb{E}\!\left[\ell(z,Y)\mid X=x\right]=\eta(-z)+(1-\eta)z=(1-2\eta)\,z.$ (79)
As in Theorem 6, $z^{\star}\in\arg\min_{z\in[0,1]}r_{\eta}(z)$ satisfies $z^{\star}=1$ if $\eta>\tfrac{1}{2}$, $z^{\star}=0$ if $\eta<\tfrac{1}{2}$, and any $z^{\star}\in[0,1]$ if $\eta=\tfrac{1}{2}$. Hence, for any $\widehat{z}\in[0,1]$,
$r_{\eta}(\widehat{z})-r_{\eta}(z^{\star})=|1-2\eta|\cdot|\widehat{z}-z^{\star}|.$ (80)
Taking expectation over $X$ gives
$R(\widehat{h}_{n})-R(h^{\star})=\mathbb{E}\!\left[\,|1-2\eta(X)|\,\big|\widehat{h}_{n}(X)-h^{\star}(X)\big|\,\right]\xrightarrow[n\to\infty]{}0.$ (81)
Let $E_{n}\coloneqq\{\widehat{C}_{n}(X)\neq C_{\rm MAP}(X)\}$. For any $\delta\in(0,1)$,
$E_{n}\subseteq\Big(E_{n}\cap\{|1-2\eta(X)|\geq\delta\}\Big)\ \cup\ \{|1-2\eta(X)|<\delta\}.$ (82)
On the event $\{|1-2\eta(X)|\geq\delta\}$, a decision mismatch (threshold $1/2$) implies $\big|\widehat{h}_{n}(X)-h^{\star}(X)\big|\geq\tfrac{1}{2}$ because $h^{\star}(X)\in\{0,1\}$ there. Therefore,
$R(\widehat{h}_{n})-R(h^{\star})$ $\geq\mathbb{E}\!\left[\,|1-2\eta(X)|\cdot\big|\widehat{h}_{n}(X)-h^{\star}(X)\big|\cdot\mathbf{1}_{E_{n}\cap\{|1-2\eta(X)|\geq\delta\}}\,\right]$
$\geq\frac{\delta}{2}\,\Pr\!\left(E_{n}\cap\{|1-2\eta(X)|\geq\delta\}\right).$ (83)
Sending $n\to\infty$ and using (81) yields
$\Pr\!\left(E_{n}\cap\{|1-2\eta(X)|\geq\delta\}\right)\xrightarrow[n\to\infty]{}0.$ (84)
Moreover, by (74), $\Pr(|1-2\eta(X)|<\delta)\to 0$ as $\delta\downarrow 0$. Combining this with (82) proves that $\Pr(E_{n})\to 0$, i.e. $\widehat{C}_{n}(X)$ agrees with $C_{\rm MAP}(X)$ with probability tending to $1$. This implies the 011 risk convergence (78). ∎